OpenStudy (anonymous):

The first side of a triangle is 8m shorter than the second side. The third side is 4 times as long as the first side. The perimeter is 26m. Find the length of each side. Could you break it down slowly? I just dont get it.

OpenStudy (anonymous):

Let second side be x m

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first side is 8m shorter than second side so (x-8) m

OpenStudy (anonymous):

and the third side is 4 times of first side so third side is = 4(x-8) m

OpenStudy (anonymous):

Now, perimeter of triangle = x + (x-8) + 4(x-8) 26 = x + x-8 +4x - 32 6x = 26 + 40 = 66 x = 11 m First Side = x-8 = 11-8 = 3 m Second Side = x = 11 m Third Side = 4(x-8) = 4*(11-8) =4*3 = 12 m

OpenStudy (anonymous):

ok thanks